# Dynamic Spiking Graph Neural Networks

### v1

@@ Introduction
Graph Neural Networks (GNNs) (Scarselli et al. 2008) have been widely applied in various fields to learn the graph representation by capturing the dependencies of nodes and edges, such as relation detection (Schlichtkrull et al. 2018; Chen, Li, and Tang 2020; Mi and Chen 2020; Xu et al. 2019b) and recommendation (Wu et al. 2022). However, most GNNs are primarily designed for static or non-temporal graphs, which cannot meet the requirement of dynamic evolution over time in practice. The established solution is to extend the dynamic graphs into sequence models directly. Typically, these methods (Yin et al. 2023b; Shi et al. 2021; Kumar, Zhang, and Leskovec 2019; Yin et al. 2022a) utilize the Recurrent Neural Networks (RNNs) (Cho et al. 2014) to capture the dynamic evolution of graphs for numerous downstream tasks such as time series prediction or graph property prediction (Yin et al. 2022b; Wieder et al. 2020; Yin et al. 2023a).
Despite the promising performance of dynamic graphs, the majority of these approaches typically involve complex structures that consume significant computational resources during training and testing. Inspired by the way the brain process information, Spiking Neural Networks (SNNs) represent the event or clock-driven signals as inference for updating the neuron nodes parameters (Brette et al. 2007). Different from traditional deep learning methods, SNNs utilize discrete spikes information instead of continuous features, resulting in significantly lower power consumption during model training. Considering the inherent characteristics of SNNs (Maass 1997; Pfeiffer and Pfeil 2018; Schliebs and Kasabov 2013), a few recent works (Zhu et al. 2022; Xu et al. 2021; Li et al. 2023) have attempted to integrate SNNs into the GNNs framework to tackle the issue of high computational complexity. These methods transform the node features into a series of spikes with Poisson rate coding, and follow a graph convolution layer with SNN neurons, which employ a membrane potential threshold to convert continuous features to spike information (Kim et al. 2020; Bu et al. 2022). Although Spiking Graph Networks (SGNs) are gradually gaining attention, the use of SNNs in dynamic graphs is still less explored, which is a more common scenario in life. To address the gap, the work focuses on the problem of spiking dynamic graph, which applies SNNs for dynamic graph node classification.
@@
$=$ $\sigma\left(\frac{1}{V_{th}}\left(\frac{\sum_{i=0}^{K-2}\lambda^{i}}{\sum_{i=0}^{K-1}\lambda^{i}}\bm{W}\bm{a}_{t+1,K-1}+\bm{F}\bar{\bm{x}}_{t+1,K-1}\right)\right)$
$-\frac{\bm{u}_{t+1,K}}{V_{th}\sum_{i=0}^{K-1}\lambda^{i}}+\frac{\bm{a}_{t,K}}{V_{th}\sum_{i=0}^{K-1}\lambda^{i}},$
@@
[…]
@@
[…]
Comparison of Runtime Complexity. We further compare the runtime complexity between Dy-SIGN with SpikeNet, JODIE, EvolveGCN, and TGAT, which is shown in Figure 2 (b). From the results, we find that the SNN-based methods are significantly more efficient than the ANNs methods. The reason for this is attributed to the fact that SNN-based methods use binary signals instead of continuous features, allowing the matrix multiplication operation to be replaced by an accumulation operation. Additionally, the proposed Dy-SIGN is slightly more efficient than SpikeNet method, the potential reason is that Dy-SIGN uses the simple form to calculate the gradient with Equation 8, ignoring the time-consuming calculation of gradient in BPTT.
@@ Conclusion
Broader Impacts and Limitations. Our research combines SNNs with dynamic graphs to address the issue of high memory consumption in dynamic graphs, as well as mitigate the problem of information loss in SNNs. This research is expected to expand a new branch in the field of GNNs: how to efficiently train and deploy GNN models in low-power and low-memory environments. The limitation of this work is that the study of combining SNNs and GNNs is still in the early stages, and how to effectively merge the two still needs to be explored. Furthermore, the method of implicit differentiation is merely one of the means, and exploring efficient solutions remains highly pressing.
@@ Appendix B Baselines
[…]
@@
[…]

### v2

@@ Introduction
Graph Neural Networks (GNNs) (Scarselli et al. 2008) have been widely applied in various fields to learn the graph representation by capturing the dependencies of nodes and edges, such as relation detection (Schlichtkrull et al. 2018; Chen, Li, and Tang 2020; Mi and Chen 2020; Xu et al. 2019b; Ju et al. 2024) and recommendation (Wu et al. 2022). However, most GNNs are primarily designed for static or non-temporal graphs, which cannot meet the requirement of dynamic evolution over time in practice. The established solution is to extend the dynamic graphs into sequence models directly. Typically, these methods (Yin et al. 2023b; Shi et al. 2021; Kumar, Zhang, and Leskovec 2019; Yin et al. 2022a) utilize the Recurrent Neural Networks (RNNs) (Cho et al. 2014) to capture the dynamic evolution of graphs for numerous downstream tasks such as time series prediction or graph property prediction (Yin et al. 2022b; Wieder et al. 2020; Yin et al. 2023a, 2024b; Pang et al. 2023).
Despite the promising performance of dynamic graphs, the majority of these approaches typically involve complex structures that consume significant computational resources during training and testing. Inspired by the way the brain process information, Spiking Neural Networks (SNNs) represent the event or clock-driven signals as inference for updating the neuron nodes parameters (Brette et al. 2007). Different from traditional deep learning methods, SNNs utilize discrete spikes information instead of continuous features, resulting in significantly lower power consumption during model training. Considering the inherent characteristics of SNNs (Maass 1997; Pfeiffer and Pfeil 2018; Schliebs and Kasabov 2013), a few recent works (Zhu et al. 2022; Xu et al. 2021; Li et al. 2023; Yin et al. 2024a) have attempted to integrate SNNs into the GNNs framework to tackle the issue of high computational complexity. These methods transform the node features into a series of spikes with Poisson rate coding, and follow a graph convolution layer with SNN neurons, which employ a membrane potential threshold to convert continuous features to spike information (Kim et al. 2020; Bu et al. 2022). Although Spiking Graph Networks (SGNs) are gradually gaining attention, the use of SNNs in dynamic graphs is still less explored, which is a more common scenario in life. To address the gap, the work focuses on the problem of spiking dynamic graph, which applies SNNs for dynamic graph node classification.
@@
$=$ $\sigma\Big{(}\frac{1}{V_{th}}\Big{(}\frac{\sum_{i=0}^{K-2}\lambda^{i}}{\sum_{i=0}^{K-1}\lambda^{i}}\bm{W}\bm{a}_{t+1,K-1}$
$+\bm{F}\bar{\bm{x}}_{t+1,K-1}\Big{)}\Big{)}-\frac{\bm{u}_{t+1,K}}{V_{th}\sum_{i=0}^{K-1}\lambda^{i}}$
$+\frac{\bm{a}_{t,K}}{V_{th}\sum_{i=0}^{K-1}\lambda^{i}},$
@@
To verify the effectiveness of the proposed Dy-SIGN, we conduct experiments on three large real-world graph datasets, i.e., DBLP (Lu et al. 2019), Tmall (Lu et al. 2019) and Patent (Hall, Jaffe, and Trajtenberg 2001). The statistics and details introduction are presented in Appendix. We compared Dy-SIGN with various competing methods, including two static graph methods ( i.e., DeepWalk (Perozzi, Al-Rfou, and Skiena 2014) and Node2Vec (Grover and Leskovec 2016)), seven dynamic graph methods (i.e., HTNE (Zuo et al. 2018), M2DNE (Lu et al. 2019), DyTriad (Zhou et al. 2018), MPNN (Panagopoulos, Nikolentzos, and Vazirgiannis 2021), JODIE (Kumar, Zhang, and Leskovec 2019), EvolveGCN (Pareja et al. 2020) and TGAT (da Xu et al. 2020)), and one spiking method SpikeNet (Li et al. 2023). The details are introduced in Appendix. As for the implementation, we follow the same settings with (Li et al. 2023) and report the Macro-F1 and Micro-F1 results under different training ratios (i.e., 40%, 60%, and 80%). Besides, we use 5% for validation. The hidden dimension of all the methods is set to 128, and the batch size to 1024. The total training epochs are 100 and the learning rate is 0.001.
@@
Comparison of Performance.
The classification results on different datasets under various training ratios are presented in Table 1 and  2. From the results, we find that the proposed Dy-SIGN achieves competitive performance compared with other methods. From the results, we have the following observations: (1) The static methods DeepWalk and Node2Vec perform worse than the others, which indicates that simply applying static graph methods to dynamic graphs ignores the contribution of historical information for representation. (2) The methods HTNE, M2DNE, DyTrid, and MPNN fail to learn meaningful representations on the large-scale Patent dataset. The potential reason behind this is the high computational complexity of these models, leading to extensive time consumption for both training and prediction. This makes it impractical to achieve competitive performance within an acceptable time frame. (3) Although the spiking methods SpikeNet and Dy-SIGN apply the binary information for representation learning, the performance is still better than JODIE, EvolveGCN, and TGAT. This phenomenon indicates that the spike-based learning method is also competitive in both computational complexity and performance compared to the traditional methods. (4) The proposed Dy-SIGN outperforms SpikeNet in half of the settings. However, as shown in Figure 2 (a), the memory consumption of Dy-SIGN is about only half of SpikeNet under the same experiment environment, demonstrating the superiority of Dy-SIGN. We attribute this to the fact that by applying the variation training method, Dy-SIGN achieves more efficient results.
Comparison of Runtime Complexity.
We further compare the runtime complexity between Dy-SIGN with SpikeNet, JODIE, EvolveGCN, and TGAT, which is shown in Figure 2 (b). From the results, we find that the SNN-based methods are significantly more efficient than the ANNs methods. The reason for this is attributed to the fact that SNN-based methods use binary signals instead of continuous features, allowing the matrix multiplication operation to be replaced by an accumulation operation. Additionally, the proposed Dy-SIGN is slightly more efficient than SpikeNet method, the potential reason is that Dy-SIGN uses the simple form to calculate the gradient with Equation 8, ignoring the time-consuming calculation of gradient in BPTT.
@@ Appendix B Baselines
Dynamic Graph Methods. We also compare the proposed Dy-SIGN with eight dynamic graph methods, i.e., HTNE (Zuo et al. 2018), M2DNE (Lu et al. 2019), DyTriad (Zhou et al. 2018), MPNN (Panagopoulos, Nikolentzos, and Vazirgiannis 2021), JODIE (Kumar, Zhang, and Leskovec 2019), EvolveGCN (Pareja et al. 2020), TGAT (da Xu et al. 2020) and SpikeNet (Li et al. 2023). We implement all the models with the codes provided by the corresponding paper. We search the hyperparameters by ranging the learning rate over {0.01,0.005,0.001, 0.0005, 0.0001} and dropout rate over {0.3,0.5,0.7} and report the best reports. For fairness, we set the embedding size of all the models to 128 as default.
@@
  • M2DNE. M2DNE (Lu et al. 2019) is a network embedding method that specifically focuses on capturing the structural and temporal properties of evolving graphs. It achieves this by incorporating a temporal attention point process and a general dynamics equation into the embedding process.